# Counting points on generic character varieties

### v2

@@ 4.1.2. Interlude on Pontryagin duality
(4) Principal series representations of $\textrm{Principal series representations of $G^{F}$}\rightarrow\mathcal{T}(G).$
@@ 4.1.6. Unipotent characters
As noted in [GM20, Remark 2.3.7], for all irreducible characters $\chi$, we have $|T^{F}|$ divides $m_{\tau}(q)$.
@@ 4.2. Evaluating the character sum $S_{\tau}$
We start by elucidating the nature of $\mathrm{Irr}(G^{F})_{\tau}$. It consists of characters $\chi_{\theta,\rho}$ as $\theta$ runs over elements of $(T^{F})^{\vee}$ satisfying $[G_{\theta}]=[L]$. Equivalently,
@@ 4.3. The auxiliary sum $\alpha_{L,S}^{\supseteq}$
Moreover, the conclusion holds if $q$ and $F$ are replaced by $q^{m}$ and $F^{m}$, for any $m\in\mathbb{Z}_{>0}$.
@@
The lemma follows from the fact (proved in [KNP23, §5.2]) that the Pontryagin dual of the embedding of finite abelian groups $(\check{T}^{F})^{W(L)}\hookrightarrow\check{T}^{F}$ is the canonical quotient map
@@
Note that the containment $[L^{F},L^{F}]\subseteq[L,L]$ may be proper (e.g. $L=G=\mathrm{PGL}_{n}$). Thus, it could happen that $S\in[L,L]\setminus[L^{F},L^{F}]$. However, this subtlety goes away after a finite base change. Indeed, there are finitely many isolated standard endoscopic groups for $G$; thus, given $S\in T^{F}$, there exists a positive integer $m$, such that $S\in[L^{F^{m}},L^{F^{m}}]$ for all isolated standard endoscopy groups. We conclude:
@@ 4.3.2. Description of $\check{T}^{W(L)}$
To proceed further, we need a description of $T^{W(L)}$. It is proved in [KNP23, §5] that, under the assumption that $G$ has connected centre, we have:
@@
Moreover, the same formula holds if $r$ is replaced by $rx$ for any positive integer $x$.
@@ 4.4. Evaluating $\alpha_{L,S}$
We now obtain an expression for $\alpha_{L,S}$ using Möbius inversion.
@@
We have seen that $\alpha_{L^{\prime},S}^{\supseteq}$ is 00 unless $L^{\prime}$ is isolated. In this case, our assumption that $G$ has connected centre implies
@@
Moreover, the equality continues to hold if $r$ is replaced by $rx$ for any positive integer $x$.
@@ 4.6. Proof of Theorem 2 in the multi-punctured case
Let ${X}$ be a character variety associated to an $n$-punctured surface group satisfying Assumption 1. Define the rational function
@@
The conclusion continues to hold if $r$ is replaced by a positive integer multiple $mr$. The same reasoning as in the single punctured case concludes the proof of Theorem 2. ∎
@@ 5.1. Palindromic property
It is known that $\mathfrak{D}$ restricts to an involution on the set of irreducible constituents of $R_{T}^{G}\,\theta$; namely, it sends the irreducible constituent corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$ to the constituent corresponding to $\epsilon\rho$, where $\epsilon$ is the sign character of $W_{\theta}$. In particular, we see that $\mathfrak{D}$ sends an irreducible character of type $[L,\rho]$ to one of type $[L,\epsilon\rho]$.
@@ 5.3.1. Genus $g=0$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
@@ 5.3.3. Genus $g>1$ or $g=1$ and non-trivial centre
We shall prove that $f_{X}(1)=0$, thus establishing Corollary 3 (c). By Remark 26, the polynomial $f_{T}(t)$ divides $m_{\tau}(t)$. This implies that if $g>1$, then $f_{T}$ divides $f_{X}$; hence $f_{X}(1)=0$. It also implies that if $g=1$, then $f_{Z}$ divides $f_{X}$. Since $Z$ is nontrivial and connected, it follows that $f_{Z}(1)=0$. Thus, $f_{X}(1)=0$.
@@ 6.1. Recollections on invariant functions
Fix a nontrivial additive character $\psi:\mathbb{F}_{q}\rightarrow\mathbb{C}^{\times}$ and a non-degenerate $G$-invariant symmetric bilinear form $\kappa:\mathfrak{g}^{F}\times\mathfrak{g}^{F}\rightarrow\mathbb{F}_{q}$. Since we have assumed that the ground characteristic is very good for $G$, such an invariant form exists, cf. [Letellier05, Proposition 2.5.12]. Note that the restriction $\kappa|_{\mathfrak{t}}$ of $\kappa$ to $\mathfrak{t}^{F}$ is also non-degenerate.
@@ 6.2. A theorem of Kazhdan and Letellier
Let $(\mathfrak{g}^{F})^{\circ}$ denote the set of split elements of $\mathfrak{g}^{F}$. The above discussions implies
@@ 6.3. Evaluating the sum $H_{\tau}$
Thus, the problem of computing $H_{\tau}$ reduces to evaluating $\beta_{L,H}$. Note that $\beta_{L,H}$ is the additive analogue of $\alpha_{L,S}$ considered in the previous section.
@@ 6.6. Multi-punctured case
Let ${Y}$ be the additive character variety associated to an $n$-punctured surface group and $(H_{1},\ldots,H_{n})\in\mathfrak{t}^{n}$ where each $H_{i}$ is regular and the tuple is generic. Then a similar analysis as in the once-punctured case gives
@@
The same argument as the previous subsection implies that $Y\times_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is polynomial count with counting polynomial $f_{Y}$. This concludes the proof of Theorem 6.
@@ 6.7. Connectedness
Our aim now is to show that $\overline{Y}$ is connected. We know that $\overline{Y}$. Thus, it suffices to show that the $E$-polynomial $f_{Y}$ is monic. The $\mathfrak{g}$-type contributing to the leading term of $f_{Y}$ is $\tau_{1}=[G,0]$. The coefficient of this term is determined by the Green polynomial $Q_{T}^{G}(1)$ because
@@ 6.8. Euler characteristic
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
@@ 7. Examples
In this section, we give examples of Euler characteristics and counting polynomials of generic character varieties. We also give tables containing $G$-types, $\mathfrak{g}$-types, and the associated data required to compute the counting polynomials. Some expressions are obtained with the help of the computer; see [Bailey] for the computer code which is based on [GHLMP96].
Throughout, $\Phi_{i}$ denotes the $i$th cyclotomic polynomial; thus, for instance
@@
7.1. Euler characteristics for $\overline{X}$ in genus 00
@@ 7.2. Euler characteristics of $\overline{X}$ when $(G,g)=(\mathrm{SO}_{5},1)$
$[\mathrm{SO}_{5}]$ $1$ $D_{8}$ $8$ $5$ $2$
$[{A_{1}\times A_{1}}]$ $1$ $S_{2}\times S_{2}$ $4$ $4$ $2$
$[{A_{1}}]$ $2$ $S_{2}$ $2$ $2$ $-4$
$[{A_{1}^{\prime}}]$ $2$ $S_{2}$ $2$ $2$ $-2$
$[T]$ $1$ $1$ $1$ $1$ $8$
@@ 7.3. Euler characteristics of $\overline{X}$ when $(G,g)=(G_{2},1)$
$[G_{2}]$ $1$ $D_{12}$ $12$ $6$ $1$
$[{A_{2}}]$ $1$ $S_{3}$ $6$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}}]$ $3$ $S_{2}\times S_{2}$ $4$ $4$ $1$
$[{A_{1}}]$ $3$ $S_{2}$ $2$ $2$ $-4$
$[{A_{1}^{\prime}}]$ $3$ $S_{2}$ $2$ $2$ $-2$
$[T]$ $1$ $1$ $1$ $1$ $12$
@@ 7.4. Euler characteristic of $\overline{{Y}}$
$\mathrm{GL}_{3}$ $1+2\cdot 3^{n-1}(3g+n-3)$ $+3^{n-2}2^{n+3}(144g^{2}+72g(n-3)+80-53n+9n^{2})$
$\mathrm{SO}_{5}$ $1+2^{n-2}+2^{2n-3}(24g+7n-22)$ $+2^{3n-5}(96g^{2}+48g(n-3)+53-35n+6n^{2})$
$G_{2}$ $\frac{1}{6}(6-2^{n}+19\cdot 3^{n+1}+5^{n})$ $+395\cdot 12^{n-2}+5g\cdot 6^{n}+6^{n-2}(48-65\cdot 2^{n})n$ $+5\cdot 3^{n}4^{n-2}(16g^{2}+8g(n-3)+n^{2})$
@@ 7.5. Types for $G=\mathrm{SO}_{5}$ and expressions for the counting polynomial $f_{X}$
$[{A_{1}\times A_{1}},2^{1}\otimes 2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},2^{1}\otimes 1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 2^{1}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 1^{2}]$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $1$ $-2$
$[T,1^{1}]$ $q^{4}\Phi_{1}^{2}$ $8^{n}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $8$
@@
$(0,3)$ $22q^{4}+12q^{3}+48q^{2}+12q+$ 2
$(0,4)$ $2q^{12}+16q^{11}+68q^{10}+208q^{9}+530q^{8}$ $+1216q^{7}+1968q^{6}+\cdots+16q+2$
$(1,1)$ $2q^{8}+4q^{7}+6q^{6}+16q^{5}+16q^{4}$ $+16q^{3}+6q^{2}+4q+2$
$(1,2)$ $2q^{16}+8q^{15}+16q^{14}+24q^{13}+32q^{12}+40q^{11}$ $+48q^{10}+24q^{9}+188q^{8}+\cdots+8q+2$
$(2,1)$ $2q^{28}+4q^{27}-4q^{25}-8q^{24}-12q^{23}+2q^{22}$ $+16q^{21}+6q^{20}+52q^{19}-40q^{18}+228q^{17}$ $-840q^{16}+708q^{15}-228q^{14}+\cdots+4q+2$
@@ 7.6. Types for $\mathfrak{so}_{5}$ and expressions for the counting polynomial $f_{Y}$
$[\mathrm{SO}_{5},1^{5}]$ 1010{10}10 $q^{4n}\Phi_{2}^{2n}\Phi_{4}^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $\Phi_{2}^{2}\Phi_{4}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},2^{2}1^{1}]$ 66{6}6 $q^{4n}\Phi_{1}\Phi_{2}^{2n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}\Phi_{2}\Phi_{4}$ $\Phi_{2}^{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}]$ 44{4}4 $\frac{1}{2}q^{4n+1}\Phi_{1}\Phi_{2}^{2}\Phi_{4}(3q+1)^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $13q+$ 1 $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}_{\star}]$ 44{4}4 $\frac{1}{2}q^{4n+1}\Phi_{1}^{2}\Phi_{2}^{n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}^{2}\Phi_{2}\Phi_{4}$ $\Phi_{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},5^{1}]$ 22{2}2 $q^{4n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $8$ $1$ $1$
$[A_{1},1^{2}]$ 44{4}4 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1},2^{1}]$ 22{2}2 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[A_{1}^{\prime},1^{2}]$ 44{4}4 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1}^{\prime},2^{1}]$ 22{2}2 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[T,0]$ 22{2}2 $3\cdot 8^{n-1}q^{4n+4}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $3$

### v3

@@ 4.1.2. Interlude on Pontryagin duality
(4) $\textrm{Principal series representations of $G^{F}$}\rightarrow\mathcal{T}(G).$
@@ 4.1.6. Unipotent characters
As noted in [GM20, Remark 2.3.7], for all irreducible characters $\chi$, we have that $|T^{F}|$ divides $m_{\tau}(q)$ (as polynomials in $q$).
@@ 4.2. Evaluating the character sum $S_{\tau}$
We start by elucidating the nature of $\mathrm{Irr}(G^{F})_{\tau}$. It consists of characters $\chi_{\theta,\rho}$ as $\theta$ runs over elements of $(T^{F})^{\vee}$ satisfying $G_{\theta}]=[L]$. Equivalently,
@@ 4.3. The auxiliary sum $\alpha_{L,S}^{\supseteq}$
This holds if $q$ and $F$ are replaced by $q^{m}$ and $F^{m}$, for any positive integer $m$.
@@
The lemma follows from the fact (proved in [KNP25, §5.2]) that the Pontryagin dual of the embedding of finite abelian groups $(\check{T}^{F})^{W(L)}\hookrightarrow\check{T}^{F}$ is the canonical quotient map
@@
Note that the containment $L^{F},L^{F}]\subseteq[L,L]$ may be proper (e.g. $L=G=\mathrm{PGL}_{n}$). Thus, it could happen that $S\in[L,L]\setminus[L^{F},L^{F}]$. However, this subtlety goes away after a finite base change. Indeed, there are finitely many isolated standard endoscopic groups for $G$; thus, given $S\in T^{F}$, there exists a positive integer $m$, such that $S\in[L^{F^{m}},L^{F^{m}}]$ for all isolated standard endoscopy groups. We conclude:
@@ 4.3.2. Description of $\check{T}^{W(L)}$
To proceed further, we need a description of $T^{W(L)}$. It is proved in [KNP25, §5] that, under the assumption that $G$ has connected centre, we have:
@@
The same formula holds if $r$ is replaced by $rx$ for any positive integer $x$.
@@ 4.4. Evaluating $\alpha_{L,S}$
We now give an expression for $\alpha_{L,S}$ using Möbius inversion.
@@
We have seen that $\alpha_{L^{\prime},S}^{\supseteq}$ is 0 unless $L^{\prime}$ is isolated. In this case, our assumption that $G$ has connected centre implies
@@
The equality continues to hold if $r$ is replaced by $rx$ for any positive integer $x$.
@@ 4.6. Proof of Theorem 2 in the multi-punctured case
Let X{X} be a character variety associated to an $n$-punctured surface group satisfying Assumption 1. Define the rational function
@@
This continues to hold if $r$ is replaced by $rm$ for any positive integer $m$. The same reasoning as in the single punctured case completes the proof of Theorem 2. ∎
@@ 5.1. Palindromic property
It is known that $\mathfrak{D}$ restricts to an involution on the set of irreducible constituents of $R_{T}^{G}\,\theta$; namely, it sends the irreducible constituent corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$ to the constituent corresponding to $\epsilon\rho$, where $\epsilon$ is the sign character of $W_{\theta}$. In particular, we see that $\mathfrak{D}$ sends an irreducible character of type $L,\rho]$ to one of type $L,\epsilon\rho]$.
@@ 5.3.1. Genus $g=0$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and $\xi$ is the rational function
@@ 5.3.3. Genus $g>1$ or $g=1$ and non-trivial centre
We shall prove that $f_{X}(1)=0$, thus establishing Corollary 3 (c). By Remark 26, the polynomial $f_{T}(t)$ divides $m_{\tau}(t)$. This implies that if $g>1$, then $f_{T}$ divides $f_{X}$; hence $f_{X}(1)=0$. It also implies that if $g=1$, then $f_{Z}$ divides $f_{X}$. Since $Z$ is non-trivial and connected, it follows that $f_{Z}(1)=0$. Thus, $f_{X}(1)=0$.
@@ 6.1. Recollections on invariant functions
Fix a non-trivial additive character $\psi:\mathbb{F}_{q}\rightarrow\mathbb{C}^{\times}$ and a non-degenerate $G$-invariant symmetric bilinear form $\kappa:\mathfrak{g}^{F}\times\mathfrak{g}^{F}\rightarrow\mathbb{F}_{q}$. Since we have assumed that the ground characteristic is very good for $G$, such an invariant form exists, cf. [Letellier05, Proposition 2.5.12]. Note that the restriction $\kappa|_{\mathfrak{t}}$ of $\kappa$ to $\mathfrak{t}^{F}$ is also non-degenerate.
@@ 6.2. A theorem of Kazhdan and Letellier
Let $(\mathfrak{g}^{F})^{\circ}$ denote the set of split elements of $\mathfrak{g}^{F}$. The above discussions imply
@@ 6.3. Evaluating the sum $H_{\tau}$
Thus, the problem of computing $H_{\tau}$ reduces to evaluating $\beta_{L,H}$. (Note that $\beta_{L,H}$ is the additive analogue of $\alpha_{L,S}$ introduced in §4.2.)
@@ 6.6. Multi-punctured case
Let Y{Y} be the additive character variety associated to an $n$-punctured surface group and $(H_{1},\ldots,H_{n})\in\mathfrak{t}^{n}$ where each $H_{i}$ is regular and the tuple is generic. Then a similar analysis as in the once-punctured case gives
@@
The same argument as the previous subsection implies $Y\times_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is polynomial count with counting polynomial $f_{Y}$. This concludes the proof of Theorem 6.
@@ 6.7. Connectedness
Our aim now is to show that $\overline{Y}$ is connected. We have shown that $\overline{Y}$ has $E$-polynomial equal to $f_{Y}$. Thus, it suffices to show that $f_{Y}$ is monic. The only $\mathfrak{g}$-type contributing to the leading term of $f_{Y}$ is $\tau_{1}=[G,0]$. The coefficient of this term is determined by the Green polynomial $Q_{T}^{G}(1)$ because
@@ 6.8. Euler characteristic
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and $\zeta$ is the rational function
@@ 7. Examples
In this section, we give examples of Euler characteristics and counting polynomials of generic character varieties. We also give tables containing $G$-types, $\mathfrak{g}$-types, and the associated data required to compute the counting polynomials. Some expressions are obtained with the help of the computer; see [Bailey] for the computer code using CHEVIE [GHLMP96].
Throughout, $\Phi_{i}$ denotes the $i$th cyclotomic polynomial; thus, for instance,
@@
7.1. Euler characteristics for $\overline{X}$ in genus 0
@@ 7.2. Euler characteristics of $\overline{X}$ when $(G,g)=(\mathrm{SO}_{5},1)$
$\mathrm{SO}_{5}]$ 11 $D_{8}$ 88 55 22
$A_{1}\times A_{1}$}] 11 $S_{2}\times S_{2}$ 44 44 22
$A_{1}$}] 22 $S_{2}$ 22 22 4-4
$A_{1}^{\prime}$}] 22 $S_{2}$ 22 22 2-2
$[T]$ 11 11 11 11 88
@@ 7.3. Euler characteristics of $\overline{X}$ when $(G,g)=(G_{2},1)$
$G_{2}]$ 11 $D_{12}$ $12$ 66 11
$A_{2}$}] 11 $S_{3}$ 66 33 22
$A_{1}\times A_{1}^{\prime}$}] 33 $S_{2}\times S_{2}$ 44 44 11
$A_{1}$}] 33 $S_{2}$ 22 22 4-4
$A_{1}^{\prime}$}] 33 $S_{2}$ 22 22 2-2
$[T]$ 11 11 11 11 $12$
@@ 7.4. Euler characteristic of $\overline{{Y}}$
$\mathrm{GL}_{3}$ $1+2\cdot 3^{n-1}(3g+n-3)$ $3^{n-2}2^{n+3}(144g^{2}+72g(n-3)+80-53n+9n^{2})$
$\mathrm{SO}_{5}$ $1+2^{n-2}+2^{2n-3}(24g+7n-22)$ $2^{3n-5}(96g^{2}+48g(n-3)+53-35n+6n^{2})$
$G_{2}$ $\frac{1}{6}(6-2^{n}+19\cdot 3^{n+1}+5^{n})$ $395\cdot 12^{n-2}+5g\cdot 6^{n}+6^{n-2}(48-65\cdot 2^{n})n +$$5\cdot 3^{n}4^{n-2}(16g^{2}+8g(n-3)+n^{2})$
@@ 7.5. Types for $G=\mathrm{SO}_{5}$ and expressions for the counting polynomial $f_{X}$
${A_{1}\times A_{1}},2^{1}\otimes 2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ 11 11 44 11 44 22
${A_{1}\times A_{1}},2^{1}\otimes 1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $^{2}\Phi_{1}^{2}\Phi_{2}^{2} qq$ 11 44 11 44 22
${A_{1}\times A_{1}},1^{2}\otimes 2^{1}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $^{2}\Phi_{1}^{2}\Phi_{2}^{2} qq$ 11 44 11 44 22
${A_{1}\times A_{1}},1^{2}\otimes 1^{2}]$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ 11 44 11 44 22
${A_{1}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $2\cdot 4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 11 22 22 22 4-4
${A_{1}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $2\cdot 4^{n}$ $\Phi_{1}^{2}\Phi_{2} qq$ 11 22 22 22 4-4
${A_{1}^{\prime}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 11 22 22 11 2-2
${A_{1}^{\prime}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $4^{n}$ $\Phi_{1}^{2}\Phi_{2} qq$ 11 22 22 11 2-2
$T,1^{1}]$ $q^{4}\Phi_{1}^{2}$ $8^{n}$ $\Phi_{1}^{2}$ 11 11 11 11 11 88
@@
$(0,3)$ $2q^{4}+12q^{3}+48q^{2}+12q+2$
$(0,4)$ $2q^{12}+16q^{11}+68q^{10}+208q^{9}+530q^{8}$ $1216q^{7}+1968q^{6}+\cdots+16q+2$
$(1,1)$ $2q^{8}+4q^{7}+6q^{6}+16q^{5}+16q^{4}$ $16q^{3}+6q^{2}+4q+2$
$(1,2)$ $2q^{16}+8q^{15}+16q^{14}+24q^{13}+32q^{12}+40q^{11}$ $48q^{10}+24q^{9}+188q^{8}+\cdots+8q+2$
$(2,1)$ $2q^{28}+4q^{27}-4q^{25}-8q^{24}-12q^{23}+2q^{22}$ $16q^{21}+6q^{20}+52q^{19}-40q^{18}+228q^{17}$ $840q^{16}+708q^{15}-228q^{14}+\cdots+4q+2$
@@ 7.6. Types for $\mathfrak{so}_{5}$ and expressions for the counting polynomial $f_{Y}$
$\mathrm{SO}_{5},1^{5}]$ 10{10} $q^{4n}\Phi_{2}^{2n}\Phi_{4}^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ 11 $\Phi_{2}^{2}\Phi_{4}$ 88 11 11
$\mathrm{SO}_{5},2^{2}1^{1}]$ 6{6} $q^{4n}\Phi_{1}\Phi_{2}^{2n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}\Phi_{2}\Phi_{4}$ $\Phi_{2}^{2}$ 88 11 11
$\mathrm{SO}_{5},3^{1}1^{2}]$ 4{4} $\frac{1}{2}q^{4n+1}\Phi_{1}\Phi_{2}^{2}\Phi_{4}(3q+1)^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $3q+1$ 88 11 11
$\mathrm{SO}_{5},3^{1}1^{2}_{\star}]$ 4{4} $\frac{1}{2}q^{4n+1}\Phi_{1}^{2}\Phi_{2}^{n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}^{2}\Phi_{2}\Phi_{4}$ $\Phi_{2}$ 88 11 11
$\mathrm{SO}_{5},5^{1}]$ 2{2} $q^{4n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ 11 88 11 11
$A_{1},1^{2}]$ 4{4} $2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 $\Phi_{2}$ 22 22 1-1
$A_{1},2^{1}]$ 2{2} $2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ 11 22 22 1-1
$A_{1}^{\prime},1^{2}]$ 4{4} $2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 $\Phi_{2}$ 22 22 1-1
$A_{1}^{\prime},2^{1}]$ 2{2} $2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ 11 22 22 1-1
$[T,0]$ 2{2} $3\cdot 8^{n-1}q^{4n+4}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}^{2}$ 11 11 11 11 33